David runs his own business. He received an offer from a company for $50,000 per year. If he sells his business, he can get an amount of $100,000. Assume the annual interest rate is 5%. David is confused to run his own business or join the company. David annual revenue from his business is enough to cover all explicit costs and has $60,000 left over. There are no additional implicit costs. Calculate David’s accounting and economic profit.
1
Expert's answer
2020-11-23T05:39:12-0500

Accounting profit = Total revenue - Explicit Cost

Economic profit = Total Revenue - ( Explicit cost + Implicit cost)

But since there are no additional implicit cost, then Economic profit is equal to Accounting profit.

Total Revenue = cost of the business+ Interest +offer

Interest =(Principal "\times" Rate "\times" "\frac{T}{100}"

Interest = 100000"\times"5"\times""\frac{1}{100}" = $ 5000

Total revenue = $100000+$50000+$5000

=$155000

Explicit cost = $155000-$60000

=$95000

Accounting profit = $155000-$95000

=$ 60000

since Accounting profit =Economic Profit

Economic Profit = $ 60000


Accounting and Economic profit = $60000